OpenStudy (anonymous):

Hillary jogged a total distance of 7 and 1 over 3miles during the months of October and November. If Hillary only jogged 1 over 6mile every day, which expression shows the number of days she went jogging? 7 and 1 over 3 multiplied by 1 over 6 7 and 1 over 3 plus 1 over 6 7 and 1 over 3 minus 1 over 6 7 and 1 over 3 divided by 1 over 6

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

anyone help plz

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Divide the total distance by the number of miles jogged per day.

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

how do u get it

OpenStudy (anonymous):

i have trouble with fractions

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

It might help to convert 7 & 1/3 into an improper fraction. Do you know how to do that?

OpenStudy (anonymous):

a little :(

jimthompson5910 (jim_thompson5910):

multiply the denominator 3 with the whole part 7 to get 21 then add on the numerator: 21+1 = 22 so \[\Large 7 \frac{1}{3} = \frac{22}{3}\] make sense?

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

now divide 22/3 over 1/6

jimthompson5910 (jim_thompson5910):

that is the same as saying 22/3 times 6/1

OpenStudy (anonymous):

ssooo what do i have to do now

jimthompson5910 (jim_thompson5910):

multiply 22/3 times 6/1

OpenStudy (anonymous):

44

jimthompson5910 (jim_thompson5910):

multiply straight across like last time

jimthompson5910 (jim_thompson5910):

44 is correct

OpenStudy (anonymous):

yes!

jimthompson5910 (jim_thompson5910):

she went jogging for 44 days

OpenStudy (anonymous):

so what is the answer these r the choices 7 and 1 over 3 multiplied by 1 over 6 7 and 1 over 3 plus 1 over 6 7 and 1 over 3 minus 1 over 6 7 and 1 over 3 divided by 1 over 6

jimthompson5910 (jim_thompson5910):

well like I said above: Divide the total distance by the number of miles jogged per day.

jimthompson5910 (jim_thompson5910):

\[\Large \text{Answer} = \frac{\text{Total Distance}}{\text{Miles jogged per day}}\]

OpenStudy (anonymous):

ohhhh